# Phase diagram of the hard-sphere potential model in three and four dimensions using a pseudo-hard-sphere potential

### v1

@@ V.2 Four dimensional equation of state
We now present the results for the four-dimensional pHS and hard-hypersphere models. Due to limitations in extending the methodology outlined in Smallenburg et al. [34] to compute the coexistence pressure for four dimensions, such results are not included in this work. Instead, we focus on presenting the equation of state obtained from BD simulations and comparing these results with EDMD simulation data and mean-field equations of state. The main findings are summarized in Fig. 4, where panel Fig. 4a) shows the results obtained for the equations of state. We use the Padé[5,4] approximation [28] derived from simulations of the four-dimensional hard-hypersphere fluid to compare with our BD simulations. We also compare with the Ivanizki equation of state [61], which applies a generalized CS approach based on the available virial coefficients for the four-dimensional hard-hypersphere fluid. Similarly, the equation of state by Amorós and Ravi follows the CS approach, but does not constrain the polynomial representation of the virial coefficients to simulation data [62]. Finally, for comparison with the solid branch, we utilize the semi-phenomenological equation of state by Speedy, which has been re-parameterized with simulation data for a $D_{4}$ lattice [63].
@@
VI Concluding remarks and perspectives
@@
In this work, we have used computer simulations and integral equations theory to systematically study the phase diagram and thermodynamic properties of the pHS model introduced by [19].
@@
First, we set out to solve the OZ equation for two specific closure relations, the Percus-Yevick and the modified Verlet approximation. We found that in both cases the PY closure is only accurate for low densities and when the density increases, the accuracy of the closure decreases considerably. On the other hand, for the three dimensional case, the MV closure relation correctly reproduced the compressibility factor of the BD simulation results, and the deviations between measurements were small, showing the high precision of the closure. For the four dimensional case, the MV closure showed less precision at higher densities, which we believe happen due to the fact that the MV closure was originally described for a three dimensional fluid. In contrast, the empirical equation of state LM [29, 50] was shown to offer highly accurate results compared to the simulation results. This equation of state was formulated to use the virial coefficients available to the date of the original work, and computer simulation data of the four-dimensional hard-sphere fluid, providing enough data to fix the free parameters in the formulation.
@@
Next, we employed BD simulations to obtain the compressibility factor of three and four dimensional fluids using the pHS model. For the particular case of the three dimensional fluid, we also simulated the solid branch of the system, obtaining a FCC crystal and computing the coexistence densities using a recent $NVT$ direct coexistence method [34]. We also computed the equation of state of the true hard sphere fluid in three and four dimensions using EDMD simulations. We compared these results to well-known mean-field equations of state, as well as other semi-phenomenological approximations, and simulations fits. We arrived at the conclusion that the pHS model can correctly reproduce the equation of state of both the three and the four dimensional fluids for all densities investigated. We also observed a good agreement with the coexistence density, and the solid branch of the three dimensional hard sphere fluid agrees with previously reported data, our own EDMD simulation data, and equations of state.
@@
A main conclusion of this work is the high accuracy of the pHS model in reproducing thermodynamic properties in three and four dimensions, something that was not tested before; in the original work only two and three dimensions were tested with much less particles that the ones employed in this contribution and within the fluid regime [19]. Higher dimensions are of interest to studies of the glass transition, and a model such as the pHS model enables the use of standard simulation techniques and continuous integration schemes, as well as optimizations well-known in the computer simulation field. Furthermore, this shows that the model can reproduce the properties of higher dimensional fluids, but the solid branches are still something that has to be studied in detail, since in higher dimensions different types of crystals are more stable than others.
@@
In closing, we have shown that the pHS model is robust enough to reproduce the thermodynamic properties of the established HS model, and with the aid of computer simulations and integral equations, we demonstrated that this model is an excellent candidate for modeling hard-like interactions in soft matter systems. It will be of interest to further test the generality of pHS in binary mixtures and polydisperse systems, since there exist mean-field approximations for the equations of state of these systems [50, 51].
@@ Acknowledgements.
The authors acknowledge financial support from Conahcyt (Grant No. CBF2023-2024-3350) and the Marcos Moshinsky Foundation.

### v2

@@ VI.2 Four dimensional equation of state
We now present the results for the four-dimensional pHS and hard-hypersphere models. Due to limitations in extending the methodology outlined in Smallenburg et al. [35] to compute the coexistence pressure for four dimensions, such results are not included in this work. Instead, we focus on presenting the equation of state obtained from BD simulations and comparing these results with EDMD simulation data and mean-field equations of state. The main findings are summarized in Fig. 5, where panel Fig. 5a) shows the results obtained for the equations of state. We use the Padé[5,4] approximation [29] derived from simulations of the four-dimensional hard-hypersphere fluid to compare with our BD simulations. We also compare with the Ivanizki equation of state [66], which applies a generalized CS approach based on the available virial coefficients for the four-dimensional hard-hypersphere fluid. Similarly, the equation of state by Amorós and Ravi (AR) follows the CS approach, but does not constrain the polynomial representation of the virial coefficients to simulation data [67]. Finally, for comparison with the solid branch, we utilize the semi-phenomenological equation of state by Speedy, which has been re-parameterized with simulation data for a $D_{4}$ lattice [68].
@@
VII Concluding remarks and perspectives
@@
In this work, we have used computer simulations and integral equation theory to systematically study the phase diagram and thermodynamic properties of the pHS model introduced by [20].
@@
First, we set out to solve the OZ equation for two specific closure relations, the Percus-Yevick and the modified Verlet approximation. We found that in both cases the PY closure is only accurate for low densities and when the density increases, the accuracy of the closure decreases considerably. On the other hand, for the three dimensional case, the MV closure relation correctly reproduced the compressibility factor of the BD simulation results, and the deviations between measurements were small, showing the high precision of the closure. For the four dimensional case, the MV closure showed less precision at higher densities, which we believe is due to the fact that the MV closure was originally described for a three dimensional fluid. In contrast, the empirical equation of state LM [30, 51] was shown to offer highly accurate results compared to the simulation results. This equation of state was formulated to use the virial coefficients available to the date of the original work, and computer simulation data of the four-dimensional HS fluid, providing enough data to fix the free parameters in the formulation.
@@
Next, we employed BD simulations to obtain the compressibility factor of three and four dimensional fluids using the pHS model. For the particular case of the three dimensional fluid, we also simulated the solid branch of the system, obtaining a FCC crystal and computing the coexistence densities using a recent $NVT$ direct coexistence method [35]. We also computed the equation of state of the true HS fluid in three and four dimensions using EDMD simulations. We compared these results to well-known mean-field equations of state, as well as other semi-phenomenological approximations, and simulations fits. We arrived at the conclusion that the pHS model can correctly reproduce the equation of state of both the three and the four dimensional fluids for all densities investigated. We also observed a good agreement with the coexistence density, and the solid branch of the three dimensional HS fluid agrees with previously reported data, our own EDMD simulation data, and equations of state.
@@
A main conclusion of this work is the high accuracy of the pHS model in reproducing thermodynamic properties in three and four dimensions, something that was not tested before; in the original work only two and three dimensions were tested with much fewer particles than the ones employed in this contribution and within the fluid regime [20]. Higher dimensions are of interest to studies of the glass transition, and a model such as the pHS model enables the use of standard simulation techniques and continuous integration schemes, as well as optimizations well-known in the computer simulation field. Furthermore, this shows that the model can reproduce the properties of higher dimensional fluids, but the solid branches are still something that has to be studied in detail, since in higher dimensions different types of crystals are more stable than others.
@@
It is also of interest to investigate the role of hydrodynamic interactions (HI) in systems of pseudo-hard-spheres and pseudo-hard-hyperspheres, particularly regarding their influence on the long-time self-diffusion coefficient. While experimental data are available for hard-sphere systems, for higher-dimensional analogues, validation is limited to comparisons with theoretical predictions or alternative simulation techniques. In this work, we have shown that the interaction potential defined in Eq. (2) can reproduce the dynamics of hard-sphere systems, closely approaching the freezing criterion predicted by Löwen et al. [60]. However, discrepancies with the experimental data are observed, which we attribute to the potential influence of HI as well as to the dated nature of the experimental results. A more systematic and comprehensive comparison between theory, simulations, and experiments appears feasible and would be valuable in assessing the extent to which the pHS model accurately captures the essential features of the dynamics and transport phenomena hard-sphere dynamics.
In closing, we have shown that the pHS model is robust enough to reproduce the thermodynamic properties of the established HS model, and with the aid of computer simulations and integral equation theory, we demonstrated that this model is an excellent candidate for modeling hard-like interactions in soft matter systems. It will be of interest to further test the generality of pHS in binary mixtures and polydisperse systems, since there exist mean-field approximations for the equations of state of these systems [51, 52].
@@ Acknowledgements.
The authors acknowledge financial support from SECIHTI (Grant No. CBF2023-2024-3350) and the Marcos Moshinsky Foundation.
Data Availability
The data that support the findings of this study are openly available in the Zenodo repository https://doi.org/10.5281/zenodo.15105653.